CocoCammron:

Solve by Substitution Method: 2x - 3y = 8 and 3x + y = 1 please help is would be a homer

REDViper:

3x+y+1=0


(1) 2x-3y+8=0
(2) From (1) Y=-1-3x
(3) Put y in (2) 2x+3(1+3x)+8=0 2x+3+9x+8=0 11x+11=0 X=-1 Put x in (3) Y=-1+3 Y=2

REDViper:

plz talk to me

Ari1234442:

That also happened to me